# 3DGSR: Implicit Surface Reconstruction with 3D Gaussian Splatting

### v1

@@ Abstract.
In this paper, we present an implicit surface reconstruction method with 3D Gaussian Splatting (3DGS), namely 3DGSR, that allows for accurate 3D reconstruction with intricate details while inheriting the high efficiency and rendering quality of 3DGS. The key insight is incorporating an implicit signed distance field (SDF) within 3D Gaussians to enable them to be aligned and jointly optimized. First, we introduce a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities. This function connects the SDF and 3D Gaussians, allowing for unified optimization and enforcing surface constraints on the 3D Gaussians. During learning, optimizing the 3D Gaussians provides supervisory signals for SDF learning, enabling the reconstruction of intricate details. However, this only provides sparse supervisory signals to the SDF at locations occupied by Gaussians, which is insufficient for learning a continuous SDF. Then, to address this limitation, we incorporate volumetric rendering and align the rendered geometric attributes (depth, normal) with those derived from 3D Gaussians. This consistency regularization introduces supervisory signals to locations not covered by discrete 3D Gaussians, effectively eliminating redundant surfaces outside the Gaussian sampling range. Our extensive experimental results demonstrate that our 3DGSR method enables high-quality 3D surface reconstruction while preserving the efficiency and rendering quality of 3DGS. Besides, our method competes favorably with leading surface reconstruction techniques while offering a more efficient learning process and much better rendering qualities. The code will be available at https://github.com/CVMI-Lab/3DGSR.
@@
Figure 1. Our method, called 3DGSR, achieves accurate 3D surface reconstruction with rich details, while maintaining the efficiency and high-quality rendering of 3DGS. The left part illustrates the performance comparison of 3DGSR with cutting-edge reconstruction methods. The right part shows the capability of our method to achieve high-quality reconstruction and rendering results simultaneously.
@@ 1. Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a new state-of-the-art approach for high-quality novel view synthesis. This method represents the geometry and appearance of a 3D scene as a set of Gaussians, which are then optimized from posed multi-view images. Thanks to its Gaussian rasterization pipeline, 3DGS achieves real-time efficiency, even when rendering high-resolution outputs. Despite its impressive rendering quality and efficiency, 3DGS generates only noisy, incomplete points for 3D geometry and struggles to accurately reconstruct the 3D surface of a scene (Tang et al., 2023a). These, however, are highly desired in various geometry-related applications, such as 3D reconstruction, geometry editing, animation, and relighting, among others.
@@
This motivates us to investigate how to enable 3DGS for high-quality surface reconstruction while preserving its rendering capabilities and speed. The inability of 3DGS to faithfully represent a 3D surface stems from its unstructured point-based geometry representation (i.e., the center of Gaussian). During optimization, the positions of these points (Gaussians) are adjusted to capture the geometry of the scene. However, the learning objective is to improve the rendering quality, which may not necessarily lead to accurate positions. The limited number of views utilized for learning further exacerbates this issue. Besides, the optimization process does not consider the geometric constraints among the points that are indeed constrained by the underlying geometry (i.e., surfaces), leading to noisy points. Further, the discrete and unstructured points make it difficult to extract 3D surfaces through post-processing, such as Poisson surface reconstruction (Kazhdan and Hoppe, 2013).
In this paper, we introduce 3DGSR, a novel approach that incorporates a neural implicit signed distance field (SDF) within Gaussians for geometry modeling, which is then jointly optimized and aligned with 3D Gaussians. First, we propose a differentiable SDF-to-opacity transformation function that converts SDF values into corresponding Gaussians’ opacities, thus connecting SDF and Gaussians. This association enforces surface constraints on Gaussians, aligns the SDF with Gaussians, and facilitates their unified optimization. Next, during learning, optimizing the Gaussians will provide supervisory signals for updating the SDF, allowing the reconstruction of intricate details. However, this only offers sparse supervisory signals to the SDF at locations occupied by the Gaussians, which proves inadequate for optimizing a continuous SDF and may lead to the formation of redundant structures. To address this issue, we further integrate volumetric rendering (Mildenhall et al., 2021; Wang et al., 2021) and align the rendered depth (normal) with that obtained from 3D Gaussians through consistency loss to optimize the SDF. This approach effectively regularizes locations not covered by 3D Gaussians, eliminating unnecessary surface reconstructions beyond the Gaussian sampling range (see Fig. 3). It is worth noting that since we only sample a few rays during each learning iteration, this does not introduce significant computational costs. The coupled representation and learning of the 3D Gaussians and the implicit SDF field form a positive feedback cycle, enabling them to improve each other through mutual learning and ultimately yield high-quality rendering and surface reconstruction. As illustrated in Fig. 1, our approach achieves accurate reconstruction and high-quality rendering results simultaneously.
Comprehensive experiments conducted on various datasets show that our method is capable of producing high-quality reconstruction results while simultaneously preserving the efficiency and qualities of 3DGS. When compared to state-of-the-art methods for surface reconstruction (Wang et al., 2021; Yariv et al., 2021; Tang et al., 2023b), our approach substantially surpasses them in rendering quality and facilitates the reconstruction of intricate structures with significantly shorter learning time.
2. Related Works
@@
Our method builds upon prior works in 3D reconstruction with multi-view images to enable fast surface reconstruction and novel view synthesis. We discuss prior works in multi-view 3D reconstruction, neural implicit representation, and point-based rendering.
Multi-view Stereo
Traditional multi-view stereo (MVS) methods have employed point-based (Barnes et al., 2009; Stereopsis, 2010; Galliani et al., 2016; Schönberger et al., 2016) and voxel-based (De Bonet and Viola, 1999; Broadhurst et al., 2001; Seitz and Dyer, 1999) representation. Point-based methods estimate the 3D positions of pixels by matching multi-view images and finding pixel correspondences. The surface can then be obtained by performing screened Poisson surface reconstruction (Kazhdan and Hoppe, 2013) on the resulting point clouds. However, these methods heavily rely on matching accuracy and struggle to reconstruct textureless regions. On the other hand, voxel-based methods cannot achieve high-quality reconstruction due to the resolution limitation caused by high memory consumption.
@@
In recent years, the fields implicitly predicted by neural networks have gained significant attention in both novel view synthesis (Lombardi et al., 2019; Sitzmann et al., 2019; Mildenhall et al., 2021) and 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Kellnhofer et al., 2021; Jiang et al., 2020; Liu et al., 2020b). Neural Radiance Field (NeRF) (Mildenhall et al., 2021) has demonstrated excellent performance in high-quality novel view synthesis by modeling scenes with a density field and a view-dependent color field. However, NeRF’s density field cannot extract high-quality surfaces. To achieve accurate surface reconstruction, NeuS (Wang et al., 2021) represents the scene with a signed distance function (SDF) and derives the alpha value from the precondition that the zero-level surface contributes the most to the volume rendering. Following NeuS, Darmon et al(Darmon et al., 2022) improves the reconstruction quality with an additional warping loss that ensures photometric consistency. Neuralangelo (Li et al., 2023b) introduces multi-resolution hash grids (Müller et al., 2022) and regularizes the model with numerical gradients and coarse-to-fine training. However, these methods often require extensive training time, which can limit their practical applicability.
Several works have focused on accelerating neural implicit representation by using implicit-explicit hybrid representation with grids (Sun et al., 2022; Karnewar et al., 2022; Müller et al., 2022; Wang et al., 2023a; Fridovich-Keil et al., 2022), trees (Yu et al., 2021; Liu et al., 2020a), points (Xu et al., 2022; Dai et al., 2023; Chen et al., 2023b), or low-rank decomposition (Chan et al., 2022; Hu et al., 2023; Chen et al., 2022). These methods speed up the training process and promote the rendering speed of radiance fields to a remarkable real-time level. However, fast and high-quality surface reconstruction from multi-view images remains a challenging problem.
Point-based Rendering
Unstructured points (Gross and Pfister, 2011) can be efficiently rendered to obtain novel-view images. However, directly rasterizing colored point clouds (Grossman and Dally, 1998) with fixed point sizes can lead to disconnected artifacts such as discontinuity, holes, and aliasing. To address these issues, previous works (Botsch et al., 2005; Pfister et al., 2000; Ren et al., 2002; Zwicker et al., 2001) extend points to larger primitives (e.g., spheres, ellipsoids, and surfels) and splat them onto the image plane. Differentiable point-based methods (Rakhimov et al., 2022; Pittaluga et al., 2019; Aliev et al., 2020) tackle these artifacts by rendering a holed latent feature map and then decoding it as a full RGB image using neural networks. Alternatively, some approaches render points as primitives (e.g., surfels and Gaussians) with trainable parameters and render them differentially (Yifan et al., 2019; Zhang et al., 2022; Kerbl et al., 2023).
@@
Most recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has demonstrated remarkable performance in terms of high-level rendering quality and real-time rendering. Subsequent works have extended 3DGS to handle 3D/4D generation (Yi et al., 2023; Tang et al., 2023a; Zielonka et al., 2023; Ren et al., 2023; Yang et al., 2023a), dynamic scenes (Luiten et al., 2023; Yang et al., 2023b, 2024; Das et al., 2023; Li et al., 2023a; Huang et al., 2023), relighting (Gao et al., 2023), and animation (Jung et al., 2023; Ye et al., 2023). Some concurrent works (Guédon and Lepetit, 2023; Chen et al., 2023a) aim to reconstruct the 3D surface from multi-view images using 3DGS. NeuSG (Chen et al., 2023a) jointly trains a NeuS (Wang et al., 2021) and 3DGS. During training, the normals derived from NeuS regularize the rotation of Gaussians, while NeuS is also regularized to ensure that the SDF values at Gaussians’ positions are close to zero. SuGaR (Guédon and Lepetit, 2023) enforces Gaussians to have limited overlap with others during the optimization of 3DGS. After that, an initial mesh can be extracted by performing Poisson reconstruction (Kazhdan and Hoppe, 2013) on Gaussians, using the shortest axis as normals. The mesh is then bound with flat Gaussians distributed on its faces and optimized to obtain refined results.
@@ 3. Preliminaries
(1) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right).$
@@
Typically, each 3D Gaussian can be treated as a conventional point cloud by using the centers of 3D Gaussians. Then, we can leverage the estimated depth (normal) and render them into a depth (normal) map using point-based alpha blending  (Gao et al., 2023) as:
(5) $\{\mathcal{D},\mathcal{N}\}=\sum_{i\in N}T_{i}\alpha_{i}\{d_{i},\mathbf{n}_{i}\}.$
Here $d_{i}$ and $\mathbf{n}_{i}$ are the depth and normal of $i$th point respectively.
@@ 4. Method
Figure 2. Pipeline of our proposed approach for implicit surface reconstruction. We model the surface with an implicit SDF field, with which the SDF value of each 3D Gaussian can be predicted. The opacity of each Gaussian is modeled as a bell-shaped function of its distance to the implicit surface. The geometry attributes of 3D Gaussians serve as a regularization for the SDF field, while the rendered image is supervised by the captured image.
Given posed multiview images $\{\mathcal{I}_{k}\}$ and a point cloud obtained from the Structure from Motion algorithm (Schönberger and Frahm, 2016), our goal is to enable 3DGS with the ability to reconstruct accurate surface while preserving its rendering quality and efficiency. To accomplish this, we propose an approach integrating neural implicit signed distance field (SDF) within 3D Gaussians as shown in Fig. 2. A hash grid and a single-layer MLP are employed to predict the SDF value at each queried location. Additionally, we introduce a differentiable SDF-to-opacity transformation function $\Phi$ to link and align SDF and 3D Gaussians. Subsequently, given a 3D Gaussian, its opacity value is determined by querying the corresponding SDF value and transforming it using the function $\Phi$. This coupled design allows the optimization of 3D Gaussians to supply supervisory signals for updating the SDF, enabling the reconstruction of intricate details. Details are unfolded in (Sec. 4.1).
However, the above is insufficient for optimizing the continuous SDF, as it can only provide supervision at locations where Gaussians are present. This results in floating artifacts in free space due to inadequate supervision (see Fig.3). To tackle this issue, we utilize volumetric rendering to generate depth and normals by querying points along the cast ray from the neural implicit SDF. Subsequently, a depth (normal) consistency loss is employed to align the rendered depth (normal) with those derived from 3D Gaussians, effectively regularizing the locations not covered by Gaussians and limiting floaters. The details are elaborated in  4.2.
4.1. Connect 3D Gaussians with Implicit SDF

### v2

@@ Abstract.
In this paper, we present an implicit surface reconstruction method with 3D Gaussian Splatting (3DGS), namely 3DGSR, that allows for accurate 3D reconstruction with intricate details while inheriting the high efficiency and rendering quality of 3DGS. The key insight is to incorporate an implicit signed distance field (SDF) within 3D Gaussians for surface modeling, and to enable the alignment and joint optimization of both SDF and 3D Gaussians. To achieve this, we design coupling strategies that align and associate the SDF with 3D Gaussians, allowing for unified optimization and enforcing surface constraints on the 3D Gaussians. With alignment, optimizing the 3D Gaussians provides supervisory signals for SDF learning, enabling the reconstruction of intricate details. However, this only offers sparse supervisory signals to the SDF at locations occupied by Gaussians, which is insufficient for learning a continuous SDF. Then, to address this limitation, we incorporate volumetric rendering and align the rendered geometric attributes (depth, normal) with that derived from 3DGS. In sum, these two designs allow SDF and 3DGS to be aligned, jointly optimized, and mutually boosted. Our extensive experimental results demonstrate that our 3DGSR enables high-quality 3D surface reconstruction while preserving the efficiency and rendering quality of 3DGS. Besides, our method competes favorably with leading surface reconstruction techniques while offering a more efficient learning process and much better rendering qualities.
@@
journal: TOGcopyright: acmlicensedjournal: TOGjournalyear: 2024journalvolume: 43journalnumber: 6article: 198publicationmonth: 12doi: 10.1145/3687952ccs: Computing methodologies Mesh geometry models
Figure 1. Our method, called 3DGSR, achieves accurate 3D surface reconstruction with rich details while maintaining the efficiency and high-quality rendering of 3DGS. The left part shows the capability of our method to achieve high-quality reconstruction and rendering results simultaneously. 2DGS (Huang et al., 2024) is the state-of-the-art Gaussian-based reconstruction method.
@@ 1. Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a new state-of-the-art approach for high-quality novel view synthesis. This method represents the geometry and appearance of a 3D scene as a set of Gaussians, which are then optimized from posed multi-view images. Thanks to its Gaussian rasterization pipeline, 3DGS achieves real-time efficiency, even when rendering high-resolution outputs. Despite its impressive rendering quality and efficiency, 3DGS generates only noisy, incomplete point clouds for 3D geometry and struggles to accurately reconstruct the 3D surface of a scene (Tang et al., 2023a). These, however, are highly desired in various geometry-related applications, such as 3D reconstruction, geometry editing, animation, and relighting, among others.
@@
This motivates us to investigate how to enable 3DGS for high-quality surface reconstruction while preserving its rendering capabilities and speed. The challenges faced by 3DGS in accurately representing a 3D surface stem from (1) its unstructured point-based geometry representation (i.e., the center of Gaussian), making it difficult to extract 3D surfaces through post-processing, such as Poisson surface reconstruction (Kazhdan and Hoppe, 2013); and (2) its elliptical shape, which is not naturally aligned with surface representations without proper constraints. Furthermore, the optimization process does not consider the geometric constraints among the points that are indeed constrained by the underlying geometry (i.e., surfaces), leading to noisy points. The limited number of views utilized for learning further exacerbates this issue.
Figure 2. Oversmooth reconstruction results from 2DGS (Huang et al., 2024) lead to poorer performance, reflected by a higher chamfer-$\mathcal{L}_{1}$ error (1.40) compared to 3DGSR (0.72).
To this end, some recent works have explored the regularization of 3DGS (Guédon and Lepetit, 2023) or the use of 2DGS (Huang et al., 2024) or Gaussian surfels (Dai et al., 2024) for learning. Subsequently, a post-processing step, employing depth fusion (Huang et al., 2024) or Poisson reconstruction (Dai et al., 2024), is required to obtain a 3D surface. Despite achieving promising results, these approaches tend to sacrifice detailed structures, resulting in smoother surfaces (Fig. 2) and suffering from limitations inherent to the reconstruction method. Moreover, these formulations essentially decouple surface reconstruction from GS learning, which not only prevents the surface from regularizing the learning of GS but also leads to alignment issues and the accumulation of errors.
In this paper, we introduce 3DGSR, a simple yet effective method that integrates a lightweight neural implicit signed distance field (SDF) within Gaussians for surface modeling, which is optimized and jointly aligned with 3D Gaussians. We start by examining the alignment of an SDF field and 3DGS, enabling unified optimization and synchronization of the surface and appearance. Inspired by the intuition that Gaussians near the surface should possess a high opacity value, we explore a tight coupling strategy that uses SDF to directly control Gaussian opacity with a differentiable SDF-to-opacity transformation function. Although this improves reconstruction quality, we find that surface reconstruction can be disrupted by high-frequency textures, particularly in low-frequency flat surface areas, leading to noisy reconstruction (Fig. 3). This issue can also be observed in the concurrent work GOF (Yu et al., 2024b) (Fig. 3), which directly leverages Gaussian opacity and volume rendering for surface extraction. Therefore, instead of tight coupling SDF and 3D Gaussians, we further investigate a loose coupling strategy that only utilizes SDF to constrain the distribution and orientations of Gaussians. This approach forces the Gaussians to be near the surface derived from SDF and aligns their orientations (the direction of the shortest axis) with the normal of the surface. By doing so, we facilitate the alignment of surface and 3D Gaussians while providing flexibility for the SDF to model surface geometry and the 3D Gaussians to capture appearance details, yielding superior results in both rendering and reconstruction (Fig. 1).
By employing the coupling strategies, we enable the optimization of Gaussians to supply supervisory signals for the SDF. However, these signals are only provided at locations occupied by the Gaussians, resulting in sparse supervision for the SDF. This sparsity proves insufficient for optimizing a continuous SDF and may lead to the creation of redundant structures. To address this issue, we integrate volumetric rendering techniques (Mildenhall et al., 2021; Wang et al., 2021) and align the rendered depth (or normal) with that obtained from the 3D Gaussians through a consistency loss. This approach effectively regularizes the regions not covered by the 3D Gaussians. It is worth noting that during each learning iteration, we only sample a few rays for rendering depth and normal using a single-layer MLP. This not only avoids introducing significant computational costs for training but also enables fast learning of SDF from the geometric cues provided by 3DGS. In turn, the improved SDF helps to regularize the 3D Gaussians effectively.
Extensive experiments conducted on multiple datasets demonstrate that our method can produce high-quality reconstruction results while maintaining the efficiency and rendering qualities of 3DGS. When compared to neural implicit reconstruction methods (Wang et al., 2021; Yariv et al., 2021; Tang et al., 2023b), our approach performs on par or even surpasses them while being significantly faster and achieving superior novel view synthesis results. In comparison to GS-based methods (Huang et al., 2024; Yu et al., 2024b), our approach outperforms them in both surface reconstruction and novel view synthesis.
Figure 3. Imperfect reconstruction from the non-uniform distributions of Gaussians. Directly defining the surface from the Gaussian points may lead the noisy reconstruction or artifacts.
2. Related Work
@@
Our method builds upon prior works in 3D reconstruction with multi-view images to enable fast surface reconstruction and novel view synthesis. We discuss close related prior works in neural implicit representation and 3D Gaussian splatting (3DGS).
@@
In recent years, the fields implicitly predicted by neural networks have gained significant attention in both novel view synthesis (Lombardi et al., 2019; Sitzmann et al., 2019; Mildenhall et al., 2021) and 3D reconstruction (Yariv et al., 2020; Niemeyer et al., 2020; Kellnhofer et al., 2021; Jiang et al., 2020; Liu et al., 2020b; Yu et al., 2024a). Neural Radiance Field (NeRF) (Mildenhall et al., 2021) has demonstrated excellent performance in high-quality novel view synthesis by modeling scenes with a density field and a view-dependent color field. However, NeRF’s density field cannot extract high-quality surfaces. To achieve accurate surface reconstruction, NeuS (Wang et al., 2021) represents the scene with a signed distance function (SDF) and derives the alpha value from the precondition that the zero-level surface contributes the most to the volume rendering. Following NeuS, Darmon et al(Darmon et al., 2022) improves the reconstruction quality with an additional warping loss that ensures photometric consistency. Neuralangelo (Li et al., 2023b) introduces multi-resolution hash grids (Müller et al., 2022) and regularizes the model with numerical gradients and coarse-to-fine training. However, these methods often require extensive training time, which can limit their practical applicability.
3D Gaussian Splatting.
@@
Most recently, 3DGS (Kerbl et al., 2023) has demonstrated remarkable performance in terms of high-level rendering quality and real-time rendering. Subsequent works have extended 3DGS to handle 3D/4D generation (Yi et al., 2023; Tang et al., 2023a; Zielonka et al., 2023; Ren et al., 2023; Yang et al., 2023a), dynamic scenes (Luiten et al., 2023; Yang et al., 2023b, 2024; Das et al., 2023; Li et al., 2023a; Huang et al., 2023), relighting (Gao et al., 2023), and animation (Jung et al., 2023; Ye et al., 2023). Some concurrent works (Guédon and Lepetit, 2023; Huang et al., 2024; Yu et al., 2024b) aim to reconstruct the 3D surface from multi-view images using 3DGS. SuGaR (Guédon and Lepetit, 2023) enforces Gaussians to have limited overlap with others during the optimization of 3DGS. After that, an initial mesh can be extracted by performing Poisson reconstruction (Kazhdan and Hoppe, 2013) on Gaussians. The mesh is then bound with flat Gaussians distributed on its faces and optimized to obtain refined results. 2DGS (Huang et al., 2024) represents the points as 2D surfels and leverages TSDF fusion to obtain the final mesh. GOF (Yu et al., 2024b) leverages ray-tracing-based volume rendering of 3D Gaussians by identifying its level. Different from the SuGaR and 2DGS, they extract the mesh from a trained tetrahedral grid with marching tetrahedra. Though these methods achieve promising reconstruction results, their surface still misses fine-grained geometry details and high-quality rendering results. Here, we derive 3DGSR combined with explicit Gaussian rasterization and implicit volumetric rendering to get detailed surface reconstruction and high-quality rendering results.
@@ 3. Preliminaries
Figure 4. Pipeline of our proposed approach for implicit surface reconstruction. We model the surface with an implicit SDF field, with which the SDF value of each 3D Gaussian can be predicted. We propose two different coupling strategies to make the distribution of Gaussians align with the implicit SDF field. The geometry attributes of 3D Gaussians serve as a regularization for the SDF field, while the rendered image is supervised by the captured image.
@@
(1) $G(x)=\exp\left(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\right),$
where $x$ is the point position in world space.
@@
Typically, each 3D Gaussian can be treated as a point by using the centers of 3D Gaussians. Then, we can leverage the estimated depth and render them into a depth map using point-based alpha blending (Gao et al., 2023) as
(5) $\mathcal{D}=\sum_{i\in N}T_{i}\alpha_{i}d_{i},$
where $d_{i}$ is the depth of $i$th point. We then project the rendering depth into 3D space and get the pseudo normal from the local plane.
@@ 4. Method
Given posed multiview images $\{\mathcal{I}_{k}\}$ and a point cloud obtained from the Structure from Motion algorithm (Schönberger and Frahm, 2016), 3DGSR integrates an implicit signed distance field (SDF) within 3D Gaussians (Fig. 4) to reconstruct high-quality surfaces while preserving the rendering quality and efficiency of 3DGS. To achieve this, 3DGSR incorporates two key designs: Alignment of Surface and 3D Gaussians (Sec. 4.1) We investiage coupling strategies to synchronize surface and 3D Gaussians, enabling their alignment and joint optimization. The tight coupling strategy utilizes a differentiable SDF-to-opacity transformation function to connect the SDF field and the opacity of 3D Gaussians; and the loose coupling strategy uses surface derived from SDF to constrain the distribution of Gaussians, make them near the surface and align the orientation of 3D Gaussians with surface normals.
By aligning 3D Gaussian and the surface, we allow unified optimization of 3D Gaussians and SDF. However, the supervisory signal only exists at locations occupied by 3D Gaussians and is thus not sufficient for supervising the dense SDF field. Thus, we propose the second design: Regularization with Volumetric Rendering (Sec. 4.2). We employ volumetric rendering to render depth and normal by querying points along the cast ray from the neural implicit SDF and align them with that derived from 3D GS with a consistency regularization loss. This provides dense supervisory signals for learning the SDF. The two designs work together to enforce surface constraints on 3D Gaussians, which in turn allows 3D Gaussians to assist in optimizing the SDF, resulting in high-quality surfaces with aligned 3D Gaussians.
Compared to previous neural implicit surface reconstruction methods (Wang et al., 2021; Yariv et al., 2021), our approach utilizes 3DGS for efficient and high-quality rendering while avoiding the dependence on time-consuming color prediction branches and volumetric rendering loss for optimization. Instead, it can be supervised by 3DGS and directly learn the SDF from optimized geometric clues (e.g., depth and normal) derived from 3DGS, making the training process more than 20 times faster (12 hours vs. 30 minutes). In comparison to Gaussian-based reconstruction methods (Huang et al., 2024), our approach unifies the optimization of 3DGS and an implicit surface, aligning them in an end-to-end manner. This eliminates the need for post-processing steps such as depth fusion or Poisson reconstruction and circumvents their limitations and error accumulations, leading to improved reconstruction quality and better alignment.
4.1. Alignment of Surfaces and Gaussians